If the denominator of a fraction is added to its numerator and the numerator is subtracted from its denominator , the fraction becomes 11/3 . If the numerator exceeds the denominator by 3 , find the fraction ?
Answer by checkley71(8403)

N=D+3 & (N+D)/(D-N)=11/3 OR SUBSTITUTING D+3 FOR N WE GET
((D+3)+D)/(D-(D+3)=11/3 OR (2D+3)/-3=11/3 OR CROSS MULTIPLYING WE GET 6D+9=-33
OR 6D=-42 OR D=-7 THEREFORE N=-7+3 OR N=-4
PROOF ((-4+(-7)))/(-7)-(-4)=11/3 OR -11/(-7+4)=11/3 OR -11/-3=11/3 OR 11/3=11/3